OpenStudy (anonymous):

2\3 x -1\4 x = 8 can someone draw out and solve the equation

OpenStudy (anonymous):

(2/3-1/4)x=8 5/12x=8 x=96/5

OpenStudy (anonymous):

how about this next one

OpenStudy (anonymous):